solve this system of equations
4x^2-5xy-y^2=10
2y=1-5x
full solution with working would be grateful
thank you

Answer by checkley77(12844)

4x^2-5xy-y^2=10
2y=1-5x
Y=(1-5X)/2
NOW REPLACE Y IN THE FIRST EQUATION WITH (1-5X)/2
4X^2-5X(1-5X)/2-[(1-5X)/2]^2=10
4X^2-(5X-25X^2)/2-(1-10X+25X^2)/4=10 THE COMMON DENOMINATOR IS 4 FOR THE LEFT SIDE.
(16X^2-10X+50X^2-1+10X-25X^2)/4=10
-41X-1=10*4
-41X=40+1
X=41/-41
X=-1 ANS.
Y=(1-5*-1)/2
Y=(1+5)/2
Y=6/2
Y=3 ANS.
PROOF:
4*-1^2-5*-1*3-3^2=10
4*1+15+-9=10
4+15-9=10
10=10